# RobotFingerPrint: Unified Gripper Coordinate Space for Multi-Gripper Grasp Synthesis

### v1

@@ Abstract
We introduce a novel representation named as the unified gripper coordinate space for grasp synthesis of multiple grippers. The space is a 2D surface of a sphere in 3D using longitude and latitude as its coordinates, and it is shared for all robotic grippers. We propose a new algorithm to map the palm surface of a gripper into the unified gripper coordinate space, and design a conditional variational autoencoder to predict the unified gripper coordinates given an input object. The predicted unified gripper coordinates establish correspondences between the gripper and the object, which can be used in an optimization problem to solve the grasp pose and the finger joints for grasp synthesis. We demonstrate that using the unified gripper coordinate space improves the success rate and diversity in the grasp synthesis of multiple grippers.  [^1] Project page is available at: https://irvlutd.github.io/RobotFingerPrint
@@ I INTRODUCTION
Grasp planning or grasp synthesis is an important problem in robotics. Traditional grasp planners such as GraspIt! [1] and OpenRAVE grasp planner [2] adopt a sampling and evaluation strategy, in which the grasps are sampled around an object and the force closure analysis [3] is utilized to evaluate the grasps sampled. These grasp planners are usually computationally expensive and may require users to specify contact points on the grippers. Recently, learning-based grasp synthesis has received more attention. These methods use generative machine learning models such as variational autoencoders [4, 5] or diffusion models [6] for grasp synthesis. These models are trained on large-scale datasets of grasps with various objects [7, 5, 8]. Therefore, they can predict grasps for unseen objects. However, most learning-based models are only trained for a specific gripper such as the Franka Panda gripper [4, 9] or a five-finger gripper [10]. Consequently, these models cannot support multiple grippers or be used for new grippers outside the training domain.
In order to train a single grasp synthesis model to support multiple grippers, researchers have explored predicting some intermediate representations of grasps, instead of directly predicting grasp pose and joint values which are specific for each gripper. In the literature, two types of intermediate representations of grasps have been explored: gripper keypoints and contact maps. For example, UniGrasp [11] and GeoMatch [12] train neural networks to predict the locations of gripper keypoints on an input object and then solve inverse kinematics to obtain the grasps. GenDexGrasp [13] predicts the contact map of a given object and then solves an optimization problem to generate grasps that comply with the contact map. These intermediate representations can be used with different grippers, even for unseen grippers outside the training domain.
@@
We notice that both gripper keypoints and contact maps have limitations in grasp synthesis. First, gripper keypoints are sparse, i.e., usually a few keypoints are sampled on the gripper. Therefore, noises at keypoint locations can significantly affect grasp quality. In GeoMatch [12], users have to manually specify these keypoints. Second, contact maps are ambiguous for grasp synthesis and their order has to pre-determined. They provide only information about the distances between the grippers and the objects, but there is no information about how each finger should touch the object.
@@
In this work, motivated by the normalized object coordinate space in object pose estimation [14] and the dense human pose estimation [15] from the computer vision community, we introduce a novel intermediate representation of grasps named the Unified Gripper Coordinate Space (UGCS). This space is the two-dimensional surface of a sphere in the three-dimensional space denoted $S^{2}$. The coordinates in this space are longitude $\lambda\in[-\pi,\pi)$ and latitude $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$. This space is shared among any robotic gripper. Given a gripper, we propose an algorithm to map the inner surface of the gripper, i.e., “palm” of the gripper, into UGCS. Therefore, every 3D point on the palm is assigned a 2D coordinate $(\lambda,\varphi)$. We consider these points as the “finger print” of the robotic gripper. In order to synthesize grasps, we train a conditional variational autoencoder (CVAE) to map a latent vector and a point cloud of an object to the unified gripper coordinate space. Specifically, CVAE predicts a unified gripper coordinate $(\lambda,\varphi)$ for every point on the surface of the object. Using the predicted unified gripper coordinates, a correspondence is established between the object surface and the gripper palm. Then we solve an optimization problem using the predicted unified gripper coordinates to obtain the grasp pose and the finger joint values of a grasp.
Compared to using gripper keypoints or contact maps for grasp synthesis, our proposed UGCS establishes dense correspondences between a gripper and an object, which is robust to noises in the CVAE predictions and resolves the ambiguities from contact maps. In addition, since the UGCS is shared among all grippers, it enables grasp transfer between grippers. We conducted grasp synthesis experiments using the MultiDex dataset [5] . Our method, RobotFingerPrint, outperforms GenDexGrasp [5] and GeoMatch [12] in terms of grasp success rate and grasp diversity.
@@ II-A Analytic Grasp Synthesis
Classic grasp planners such as GraspIt! [1] and OpenRAVE grasp planner [2] use analytic approaches to access grasp quality. Such methods employ task wrench space analysis [16, 17] or force closure analysis [3, 18] to assess the feasibility of a proposed grasp. These traditional grasp synthesis methods usually use sampling-based optimization algorithms, such as simulated annealing to sample grasps, and then evaluate them, which makes the optimization quite slow. Some methods also require the user to manually specify the preferred contact regions on the gripper surface prior to starting the grasp generation process. Recent analytic grasp synthesis leverages Differentiable Force Closure (DFC) [19] in grasp optimization, where the optimization runtime is significantly improved. The main limitation of analytic methods is that the success rate of grasping may not be guaranteed. Because force closure analysis cannot account for other factors in grasping such as friction, materials of objects, noises in control, etc.
@@ II-B Learning-based Grasp Synthesis
To improve speed and success in grasp synthesis, more efforts have recently been devoted to learning-based grasp synthesis. Datasets of grasps are first constructed for learning. These grasps are then verified as successful in a physics simulator. Representative datasets include Acronym [7], UniGrasp  [11], DexGraspNet [20], Fast-Grasp’D [8], GenDexGrasp [13] and MultiGripperGrasp [21]. These datasets contain different numbers of grippers, objects, and grasps, which can be used to learn generative models for grasp synthesis such as variational auto-encoders or diffusion models.
@@
Grasp synthesis for a single gripper. Most learning-based methods can only plan grasps for a specific type of robotic gripper. For example, [4, 22, 9] deal with parallel jaw grippers while FFHNet [10] is designed for a five-finger gripper. Unlike analytic approaches, these learning-based methods cannot plan grasps for a variety of robotic grippers.
@@
Grasp synthesis for multiple grippers. Recently, a few approaches are proposed for grasp synthesis of multiple grippers. AdaGrasp [23] uses gripper features to convolve with scene features to generate grasps. UniGrasp [11] trains a neural network to predict finger-tip contact points on objects and then solves inverse kinematics to generate grasps. However, both AdaGrasp and UniGrasp cannot scale-up to grippers with more than 3 fingers. GenDexGrasp [13] trains a conditional variational autoencoder to predict contact map of a given object and then solves an optimization problem using the predicted contact map to generate a grasp. GeoMatch [12] follows a similar idea as UniGrasp [11] to predict keypoints of grippers on object surfaces and then solve inverse kinematics to compute grasps. Different from existing methods, we introduce a novel contact map representation that encodes dense correspondences of gripper surface and object surface. This representation improves success rate and diversity in multi-gripper grasp synthesis across several scenarios.
III Method
III-A Unified Gripper Coordinate Space (UGCS)
The Unified Gripper Coordinate Space (UGCS) is a coordinate space that represents the points on the interior surface of a gripper by mapping them onto the surface of a sphere. This coordinate space is common to different grippers and provides a dense representation of the interior surface. A sphere was chosen because of its smooth surface point representation and the fact that similar contact regions can be marked unambiguously across grippers with varying scales. The UGCS representation assigns a normalized coordinate $(\lambda,\varphi)$, with each component in $[0,1]$ obtained after normalizing the longitude and latitude in their ranges $\lambda\in[-\pi,\pi)$ and $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$. Next, we present the method for constructing such a unified coordinate space across different grippers via their grasps on an associated maximal sphere.
@@ III-A Unified Gripper Coordinate Space (UGCS)
Figure 2: Max spheres across all the 12 different grippers including the gripper set from MultiGripperGrasp [21] dataset.
@@
Figure 3: Maximal Graspable Sphere: We test for multiple spheres with varying radii in parallel via Isaac Sim [24]
III-A1 Maximal Sphere Computation
@@
Since the geometry and kinematics amongst grippers differ significantly, there is no single sphere that can be used to assign this coordinate mapping. Hence, we first created a test bed in simulation to obtain the maximal graspable sphere for each gripper, and used it to assign the coordinate mapping. We used Nvidia’s Isaac Sim [24] robotic simulator based on the MultiGripperGrasp [21] toolkit. The simulation consisted of top-down grasps of spheres with increasing radii. These spheres were initialized with their pole touching the gripper palm, and the position, joint values of the gripper at the moment of contact were recorded. Each grasp is simulated for 3 seconds after gripper comes in contact with the object. A grasp is considered successful if, at the end of the grasp the sphere is still in contact with the gripper. Fig. 3 shows the process of sphere-grasping tests in Isaac Sim. The maximal sphere is then determined by the maximum radius within the subset of successful grasps. The maximal graspable sphere of all grippers in the MultiGripperGrasp [21] and MultiDex [5] datasets can be seen in Fig. 2.
III-A2 UGCS for Gripper Points
@@
We use the maximal sphere for each gripper to assign a $(\lambda,\varphi)$ coordinate to the gripper’s interior surface points. We only consider the interior surface points as they are in contact or close proximity with object for any grasp. This also reduces the ambiguity in which gripper regions are supposed to contact the object, leading to stable grasps. The interior gripper surface points set $P_{G}$ is computed by uniformly sampling a set $R$ of unit ray directions from the sphere center and checking for ray-mesh intersections $\mathcal{I}_{1}(\hat{r},G)$ with the gripper’s meshes.
$P_{G}=\{v_{g}\in G\;|\;\exists\;\hat{r}\in R,\;\text{s.t}\;\mathcal{I}_{1}(\hat{r},G)=v_{g}\}.$
Each intersecting unit ray direction $\hat{r}$ can alternatively be viewed as surface points on a unit sphere with some spherical coordinate $(radius=1;\lambda,\varphi)$. If the unit ray $\hat{r}_{g}$ was used to obtain an interior surface point $v_{g}\in P_{G}$, then we simply assign the ray’s angular component of the spherical coordinate $(\lambda_{\hat{r}_{g}},\varphi_{\hat{r}_{g}})$ as the gripper point’s coordinate: $(\lambda_{v_{g}},\varphi_{v_{g}}):=(\lambda_{\hat{r}_{g}},\varphi_{\hat{r}_{g}})$. Consequently, with a uniform set of rays, we get dense gripper surface points and their associated coordinates. For $M$ gripper interior surface points, we obtain a UGCS gripper representation with shape $(M,2)$ and label this as $\Phi_{G}$. Note that unlike manually selected sparse gripper key points, the UGCS representation is not affected by the ordering of points or the number $M$. Fig. 4 illustrates the idea behind obtaining $\Phi_{G}$ for each gripper.
III-B Grasp Representation
Given a set of grasps on an object with a given gripper, we show how to utilize the UGCS representation with a coordinate map $\Phi$ on the object point cloud. Note that the gripper point cloud $P_{G}$ with its $M$ interior surface points and the corresponding UCGS coordinate map $\Phi_{G}$ are fixed across the different grasps on the object set. Consider an object point cloud $O$ having $N$ points and a grasp for gripper $G$ with its fixed interior surface point set $P_{G}$ in the object frame. Now, we have $N$ object points and $M$ gripper points in the object frame, with each gripper point having a corresponding $(\lambda,\varphi)$ coordinate. For the object’s coordinate map $\Phi_{O}$, we assign coordinates from the closest gripper point.
$\Phi_{O}^{(v_{o})}=\Phi_{G}^{(v_{g}^{*})}\;\text{where}\;v_{g}^{*}=\operatorname*{\arg\!\min}_{v_{g}\in P_{G}}||v_{o}-v_{g}||.$
In this way, we obtain a UGCS coordinate for each object point in $O$. The object map $\Phi_{O}$ provides an alternate grasp representation on the object, and its object-centric form is not constrained by the differences in grippers. A drawback in the above formulation is that it assigns a coordinate $(\lambda,\varphi)$ to each object point $v_{o}$ regardless of its distance to the gripper and without considering whether it is actually in contact or not. In order to represent such no-contact object points in $P_{O}$, we simply map them to a default coordinate of $(\lambda=0,\varphi=0)$.
Object points with such a coordinate give signal to the learning model that they are not in contact with the object. We can safely assign this specific coordinate as we are guaranteed that no gripper point ever gets assigned this coordinate. This specific coordinate represents the opposite pole of the sphere from where it was grasped. The maximal sphere simulation test ensures that no gripper point gets projected to this coordinate (see Fig. 4). Finally, similar to a contact-map dataset on multi-gripper grasps, we create a coordinate-map dataset with grasps on the objects. Fig. 5 shows some examples of such grasp coordinate maps on object surface with the no-contact regions being transparent.
@@ III-B Grasp Representation
III-C Model Architecture

### v2

@@ Abstract
We introduce a novel grasp representation named the Unified Gripper Coordinate Space (UGCS) for grasp synthesis and grasp transfer. Our representation leverages spherical coordinates to create a shared coordinate space across different robot grippers, enabling it to synthesize and transfer grasps for both novel objects and previously unseen grippers. The strength of this representation lies in the ability to map palm and fingers of a gripper and the unified coordinate space. Grasp synthesis is formulated as predicting the unified spherical coordinates on object surface points via a conditional variational autoencoder. The predicted unified gripper coordinates establish exact correspondences between the gripper and object points, which is used to optimize grasp pose and joint values. Grasp transfer is facilitated through the point-to-point correspondence between any two (potentially unseen) grippers and solved via a similar optimization. Extensive simulation and real-world experiments showcase the efficacy of the unified grasp representation for grasp synthesis in generating stable and diverse grasps. Similarly, we showcase real-world grasp transfer from human demonstrations across different objects. [^1] Project page with code and videos is available at: https://irvlutd.github.io/RobotFingerPrint
@@ I INTRODUCTION
Grasp synthesis is a fundamental problem in robot manipulation. Traditional grasp planners such as GraspIt! [1] and OpenRAVE [2] adopt a sampling and evaluation strategy, in which the grasps are sampled around an object and a force closure analysis [3] is used to evaluate the grasps. These grasp planners are usually computationally expensive and may require users to manually specify contact points on the grippers. Recently, learning-based grasp synthesis has received more attention for generalization to objects and grippers. These methods use generative machine learning models such as variational autoencoders [4, 5] or diffusion models [6] for grasp generation. They are trained on large-scale grasp datasets with various objects, which enable them to predict grasps for unseen objects [7, 5, 8]. However, most learning-based models are trained only for a specific gripper [4, 9, 10]. Consequently, these models lack the ability to synthesize grasps for robotic hands outside their training domain.
In order to train a single grasp synthesis model to support multiple grippers, researchers have explored predicting intermediate grasp representations instead of directly predicting the grasp pose and joint values which are specific for each gripper. In the literature, two types of intermediate grasp representation have been developed: gripper keypoints and contact maps. For example, UniGrasp [11] and GeoMatch [12] train neural networks to predict the locations of gripper keypoints on an input object and then solve inverse kinematics to obtain the grasps. Furthermore, GenDexGrasp [5] predicts a contact map of a given object and then solves an optimization problem to generate grasps that comply with the generated contact map. These intermediate grasps representations have been used to synthesize grasps for various grippers, even for unseen grippers outside the training domain.
@@
We illustrate the limitations in both gripper keypoints and contact maps for grasp synthesis in Fig. 1. On one hand, gripper keypoints are sparse. Therefore, noise for any predicted keypoint location can significantly affect grasp quality. Users must manually specify these keypoints [12], and the keypoint prediction order is predetermined, thus inhibiting flexibility. Similarly, contact maps are ambiguous for grasp synthesis when mapping grippers to contact locations. They only provide information about the distances between the gripper and the object, but there is no information on how and where the gripper fingers should touch the object [5]. Consequently, these representations are neither suitable for transferring one grasp from one gripper to another gripper.
@@
Motivated by the normalized object coordinate space in object pose estimation [13] and dense human pose estimation [14], we introduce a novel intermediate grasp representation named the Unified Gripper Coordinate Space (UGCS). This space, denoted as $S^{2}$, is the two-dimensional surface of a sphere in a three-dimensional space. Any point within the UGCS is defined by its longitude $\lambda\in[-\pi,\pi)$ and its latitude $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$. This space is intended to be shared between all robotic hands and used to map the inner surface of the grippers (gripper “palm”). Here, every 3D point on the palm is assigned a 2D coordinate $(\lambda,\varphi)$. We refer to these points as the “finger print” of the robotic gripper. For grasp synthesis, we train a Conditional Variational Auto-Encoder (CVAE) to map a latent vector and a point cloud of an object to the UGCS. Specifically, CVAE predicts the UGCS points $(\lambda,\varphi)$ for every point on the surface of the object. Then, an objective function is optimized using the predicted coordinates to obtain the synthesized grasp configuration. Cross-gripper grasp transfer is computed by first initializing a mutual correspondence between surface points for source and target grippers, via their respective coordinate maps. Then the target gripper grasp is inferred from an optimization that tries to minimize the distance between the corresponding points.
Compared to keypoints and contact maps as grasp representations, our proposed UGCS establishes dense correspondences between a gripper and an object. We conducted grasp synthesis experiments using the MultiDex dataset [5] and compared it to methods that use alternative intermediate grasp representations [5, 12]. Our method, RobotFingerPrint (RFP), was found to compare favorably against prior works in terms of both grasp success rate and grasp diversity. Furthermore, we tested our method with out-of-domain Fetch robot [15] gripper by performing real-world manipulation on the SceneReplica [16] benchmark. The grasps generated by the RFP were able to outperform on the model-based grasp scenario with the GraspIt [1] baseline of the benchmark. We also showed the effectiveness in grasp transfer from human hand to the Fetch gripper. Such a transfer pipeline can be an effective tool in learning from demonstration scenarios with purely visual information. Our contributions are summarized as follows.
  • We introduce the unified gripper coordinate space that establishes dense correspondences across different grippers via a novel shared representation space.
  • Diverse and stable object-centric grasp synthesis across in/out-domain grippers and object sets with stronger pose initialization.
  • Generalized grasp transfer across different grippers without requiring manual re-targeting.
@@ II-A Analytic Grasp Synthesis
Classical grasp planners such as GraspIt! [1] and OpenRAVE grasp planner [2] use analytic approaches to access grasp quality. Such methods employ task wrench space analysis [17, 18] or force closure analysis [3, 19] to assess the feasibility of a proposed grasp. These traditional grasp synthesis methods usually use sampling-based optimization algorithms, such as simulated annealing to sample grasps, and then evaluate them, which makes the optimization quite slow. Some methods also require the user to manually specify the preferred contact regions on the gripper surface prior to starting the grasp generation process. Recent analytic grasp synthesis leverages Differentiable Force Closure (DFC) [20] in grasp optimization, where the optimization runtime is significantly improved. The main limitation of analytic methods is that the success rate of grasping may not be guaranteed because force closure analysis cannot account for other factors in grasping such as friction, materials of objects and noises in control.
@@ II-B Learning-based Grasp Synthesis
To improve speed and success in grasp synthesis, recent efforts have been devoted to learning-based grasp synthesis. Datasets have been designed for machine learning methods on grasping via grasp stability verification in physics simulators. Representative datasets include Acronym [7], UniGrasp [11], DexGraspNet [21], Fast-Grasp’D [8], GenDexGrasp [5] and MultiGripperGrasp [22]. These datasets contain different numbers of grippers, objects, and grasps, which can be used to learn generative models for grasp synthesis such as variational auto-encoders or diffusion models.
II-B1 Single-Gripper Models
@@
Most learning-based methods can only plan grasps for a specific type of robotic gripper. For example, [4, 23, 9] deal with parallel jaw grippers while FFHNet [10] is designed for a five-finger gripper. Unlike classical approaches, these learning-based methods cannot plan grasps for a variety of robotic hands.
II-B2 Multi-Gripper Models
@@
In recent years, a few approaches have been proposed for grasp synthesis of multiple grippers. For example, AdaGrasp [24] convolves over both normalized gripper features and scene features to predict grasps, which in turn enables it to generalize over robotic hands. UniGrasp [11] trains a neural network to predict finger-tip contact points on objects, and then solves inverse kinematics to generate grasps. However, both AdaGrasp and UniGrasp cannot scale-up to grippers with more than 3 fingers. On the other hand, GenDexGrasp [5] use a CVAE to predict contact maps on a given object and then solve an optimization problem using the predicted contact map to generate a grasp. GeoMatch [12] follows a similar idea as UniGrasp [11] to predict keypoints of grippers on object surfaces and then solve inverse kinematics to compute grasps. We introduce a new method for multi-gripper grasp synthesis based on our UGCS representation.
II-C Grasp Transfer
For grasp transfer, previous work either focuses on the problem of cross-object transfer [25, 26] or formulates it as retrieval from an offline dataset [27] that inhibits generalization to unseen objects and grippers. Unlike existing methods, we introduce a novel grasp representation that encodes dense correspondences of gripper surfaces and object surfaces. This representation enables grasp transfer that generalizes to novel grippers and objects.
@@ II-C Grasp Transfer
Figure 2: Max spheres across all the 12 different grippers including the gripper set from the MultiGripperGrasp [22] dataset.
III Unified Gripper Coordinate Space
The Unified Gripper Coordinate Space (UGCS) is a coordinate space that maps the points on the interior surface of a gripper onto the surface of a sphere. This coordinate space is common to different grippers and provides a dense representation of the interior surface. A sphere was chosen because of its smooth surface point representation and the fact that similar contact regions can be marked unambiguously across grippers with varying scales. The UGCS representation assigns a normalized coordinate $(\lambda,\varphi)$, with each component in $[0,1]$ obtained after normalizing the longitude and latitude in their ranges $\lambda\in[-\pi,\pi)$ and $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$. Next, we present the method for constructing such a unified coordinate space across any gripper via their grasp on an associated maximal sphere.
@@ III Unified Gripper Coordinate Space
Figure 3: Maximal Graspable Sphere: We test for multiple spheres with varying radii in parallel using Isaac Sim [28]
III-A Maximal Sphere Computation
@@
Since the geometry and kinematics between grippers differ significantly, there is no single sphere that can be used to assign this coordinate mapping. Hence, we created a test bed in simulation to obtain the maximal graspable sphere of each gripper for assigning the coordinate mapping. We used the NVIDIA Isaac Sim [28] robotic simulator where the simulation consisted of top-down grasps of spheres with increasing radii. These spheres were initialized with a common pole position and each gripper performed an encompassing grasp from the south pole. Each grasp was simulated for 3 seconds after the gripper came in contact with the sphere. The maximal spheres of grippers in the MultiGripperGrasp [22] and MultiDex [5] datasets can be seen in Fig. 2, while Fig. 3 shows the process of sphere-grasping tests in Isaac Sim.
III-B UGCS for Gripper Points
@@
We use the maximal sphere for each gripper to assign a $(\lambda,\varphi)$ coordinate to the interior surface points of the gripper. Interior surface points are considered as they are close to the objects. This also reduces the ambiguity in which gripper points are supposed to contact the object, leading to stable grasps. The interior gripper surface points set $P_{G}$ is computed by uniformly sampling a set $R$ of unit ray directions from the sphere center and checking for ray-mesh intersection $\mathcal{I}_{1}(\hat{\mathbf{r}},G)$ with the mesh $G$ of the gripper:
$P_{G}=\{\mathbf{v}_{g}\in G\;|\;\exists\;\hat{\mathbf{r}}\in R,\;\text{s.t.,}\;\mathcal{I}_{1}(\hat{\mathbf{r}},G)=\mathbf{v}_{g}\},$ (1)
Here $\mathbf{v}_{g}$ is the intersection point of the ray $\hat{\mathbf{r}}$ and the gripper mesh $G$. Each intersecting unit ray direction $\hat{\mathbf{r}}$ is also viewed as a surface point on a unit sphere with spherical coordinate $(\lambda,\varphi)$. We then assign the ray’s angular component of the spherical coordinate as the gripper point’s coordinate: $(\lambda_{\mathbf{v}_{g}},\varphi_{\mathbf{v}_{g}}):=(\lambda_{\hat{\mathbf{r}}},\varphi_{\hat{\mathbf{r}}})$. With a uniform set of rays, we obtain dense gripper surface points and their associated coordinates. For $M$ gripper points, we obtain the UGCS representation with shape $(M,2)$ and denote this as $\Phi_{G}$. Note that unlike manually selected sparse gripper key points, the UGCS representation is not affected by the ordering of points or the number $M$. Fig. 4 illustrates the idea behind obtaining $\Phi_{G}$ for each gripper. The gripper point cloud $P_{G}\;;|P_{G}|=M$ and its UGCS coordinate map $\Phi_{G}$ are fixed for the gripper across its different grasps on objects.
@@
III-C Grasp Representation
In this section, we detail how to represent grasps in an object-centric manner using the UGCS coordinate map $\Phi$ on the object’s point cloud. Consider an object $O$ ’s point cloud $P_{O}$ and a grasp from gripper $G$ with its fixed surface point set $P_{G}$ and coordinates $(\lambda,\varphi)$ for each gripper point. To construct the coordinate map $\Phi_{O}$ of the object, we assign coordinates from the closest gripper point:
$\Phi_{O}^{(\mathbf{v}_{o})}=\Phi_{G}^{(\mathbf{v}_{g}^{*})},\;\text{where}\;\mathbf{v}_{g}^{*}=\operatorname*{\arg\!\min}_{\mathbf{v}_{g}\in P_{G}}||\mathbf{v}_{o}-\mathbf{v}_{g}||,$ (2)
Here $\mathbf{v}_{o}$ indicates an object point. In this way, we obtain the UGCS coordinate for each object point. The object map $\Phi_{O}$ provides an alternate grasp representation on the object, and its object-centric form is not constrained by the differences in grippers. For non-contact object points in $P_{O}$ (determined by a threshold of 1cm on gripper-object distance), we simply map them to a default coordinate of $(\lambda=0,\varphi=0)$. This is the sphere’s north pole with original coordinates as $(-\pi/2,-\pi)$, and is never mapped to a gripper point as a consequence of maximal sphere test (see Fig. 4). Finally, similar to a contact-map dataset on multi-gripper grasps, we create a coordinate-map dataset with grasps on the objects. Fig. 5 shows some examples of such grasp coordinate maps on object surface with the no-contact regions being transparent.
IV Grasp Synthesis
IV-A Model Architecture